Theory and Applications of Categories, Vol. 32, No. 46, 2017, pp. 1588–1600.

A CHARACTERIZATION OF THE HUQ COMMUTATOR

VAINO TUHAFENI SHAUMBWA

Abstract. We study a categorical commutator, introduced by Huq, defined for a pair of coterminal morphisms. We show that in a normal unital category ℂ with finite colimits, the normal closure of the regular image of the Huq commutator of a pair of subobjects under an arbitrary morphism is the same as the Huq commutator of their respective regular images. Then we use this property to characterize the Huq commutator as the largest commutator satisfying certain properties

1. Introduction

The concept of the commutator of subgroups of a given group has been generalized in various ways from groups to various categorical contexts. In the present paper, we study the concept of commutator in the sense of [Huq, 1986], which is derived from the notion of commuting morphisms introduced by [Huq, 1986] and developed further by several other mathematicians working in categorical algebra. The commutator introduced by [Smith, 1976] for congruences in a Mal’tsev variety and later on defined categorically for internal equivalence relations by [Pedicchio, 1995], had been characterized by [Hagemann, Herrmann, 1979] as the largest commutator defined on congruence lattices satisfying certain properties. In the present paper, we characterize the Huq commutator as the largest commutator defined for subobjects satisfying certain properties; a result that could be thought of as a Huq commutator version of Herrmann and Hagemann characterization of Smith commutator. In Section 2 we recall the notion of commuting morphisms, as well as the construction of the Huq commutator of a pair of coterminal morphisms in a normal unital category with finite colimits. Furthermore, we recall (in a slightly weaker context) the categorical definition of the Higgins commutator as formulated by [Mantovani, Metere, 2010] for a pair of subobjects in an ideal-determined category [G. Janelidze et al., 2010]. Relating these two commutators (Huq and Higgins), we prove (in Proposition 2.2) that in a normal unital category ℂ with finite colimits, the Huq commutator is the normal closure of the Higgins commutator. This result was initially proven for an ideal-determined unital category in [Mantovani, Metere, 2010, Proposition 5.7].

The author is very grateful to James Gray for his support. Valuable suggestions made by the referee are highly appreciated. The author acknowledges the financial assistance by DAAD and NRF.

Received by the editors 2016-04-13 and, in final form, 2017-12-19.

Transmitted by Walter Tholen. Published on 2017-12-31.

2010 Mathematics Subject Classification: 18B30, 18A05, 18B35, 18A32 .

Key words and phrases: Commuting morphisms; Commutator; Normal, Ideal-determined, Unital category.

\begin{matrix}c \\ ◯ \end{matrix} Vaino Tuhafeni Shaumbwa, 2017. Permission to copy for private use granted.

In Section 3 we show (in Theorem 3.2) that in a normal unital category ℂ with finite colimits, the normal closure of the regular image of the Huq commutator of a pair of subobjects under an arbitrary morphism is the same as the Huq commutator of their respective regular images. In the last section we prove (in Theorem 4.5) that in a normal unital category ℂ with finite colimits, the Huq commutator is the largest commutator defined for subobjects satisfying certain properties.

2. Preliminaries

Recall that a unital category ℂ [Bourn, 1996] is a pointed finitely complete category with the property that for each pair of objects X and Y, the morphisms ⟨1, 0⟩ : X ⟶ X × Y and ⟨0, 1⟩ : Y ⟶ X × Y are jointly strongly-epimorphic. Following [Z. Janelidze, 2010], we define a normal category to be a pointed regular category where every regular epimorphism is a normal epimorphism. In this paper we will call a subobject h : H ↣ X normal if h is the kernel of some morphism.

Commuting morphisms [Huq, 1986] A pair of coterminal morphisms f : X ⟶ Y and g : Z ⟶ Y in a unital category ℂ commutes when there exists a (necessarily unique) morphism φ : X × Z → Y making the diagram

X Y X × Z Z f ⟨1, 0⟩ g ⟨0, 1⟩ φ

commute. Following [Bourn, 2002], we will call φ the cooperator of f and g. Note that in the category of groups, a pair of inclusions of subgroups H ↪ G and K ↪ G commutes in the above sense if and only if the subgroups commute element-wise in G. In a regular unital category ℂ, given a diagram as

X′ X Z Y Y′, \begin{matrix}f \\ g \end{matrix} \begin{matrix}f′ \\ g′ \end{matrix}

where f′ and g′ are regular epimorphisms, f ∘ f′ and g ∘ g′ commute if and only if f and g commute (see [Borceux, Bourn, 2004, Proposition 1.6.4]). For this reason, we will only define the Huq commutator for subobjects. In this paper, for a morphism t : V ⟶ W in a regular category, by “regular epimorphismmonomorphism factorization of t” we will mean the factorization of t as a regular epimorphism followed by a monomorphism.

For a pair of subobjects h : H ↣ X and k : K ↣ X of X in a normal unital category ℂ with finite colimits, we will denote by κ_{H,K} : H ⋄ K ↣ H + K the kernel of the canonical morphism

[ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] : H + K ⟶ H × K

(which is a normal epimorphism in a normal unital category with finite colimits). The object H ⋄ K is the co-smash product of H and K. Huq commutator [Huq, 1986] For a pair of subobjects h : H ↣ X and k : K ↣ X of X in a normal unital category ℂ with finite colimits, the Huq commutator [H, K]_{Q} of h : H ↣ X and k : K ↣ X is the smallest normal subobject of X for which the composites q ∘ k and q ∘ h in the diagram

H ; X ; X/[H, K]_{Q} ; X ; K ; \begin{matrix}h \\ k \end{matrix} ; \begin{matrix}q \\ q \end{matrix}

where q is the cokernel of [H, K]_{Q} ↣ X, commute. The Huq commutator of a pair of subobjects h : H ↣ X and k : K ↣ X of X in a normal unital category ℂ with finite colimits always exists, and can be constructed as the kernel of q in the diagram

H X H × K K Q [H, K]_{Q} \begin{matrix}h \\ k \end{matrix} \begin{matrix}⟨1, 0⟩ \\ ⟨0, 1⟩ \end{matrix} \begin{matrix}ker(q) \\ φ \end{matrix} q

where Q is the colimit of the solid arrows (see [Bourn, 2004]). Or equivalently (see e.g [Mantovani, Metere, 2010, Proposition 5.5]) as the kernel of q in the following pushout

H + K Q. X ⌜ H × K [H, K]_{Q} \begin{matrix}φ \\ [h, k] \end{matrix} ker(q) \begin{matrix}[ & 1 0 0 1 & ] \\ q \end{matrix}

The morphism q is the universal arrow for which the composites q ∘ h and q ∘ k commute, and is clearly a normal epimorphism (being the pushout of a normal epimorphism) in a normal unital category with finite colimits.

Higgins commutator [Higgins, 1956] [Mantovani, Metere, 2010] Let h : H ↣ X and k : K ↣ X be a pair of subobjects of X in a normal unital category ℂ with finite colimits. The Higgins commutator [H, K]_{H} of h : H ↣ X and k : K ↣ X is obtained as the regular image of the kernel κ_{H,K} : H ⋄ K ↣ H + K under the morphism [h, k] : H + K ⟶ X (see the diagram below)

\begin{matrix}H + K \\ [H, K]_{H} \\ X. \\ H ⋄ K \\ [h, k] \\ κ_{H,K} \end{matrix}

Ideal-determined categories [G. Janelidze et al., 2010] A category ℂ with finite colimits is called ideal-determined if it is normal and regular images of kernels under regular epimorphisms are kernels. For a subobject h : H ↣ X of X in a normal category ℂ with finite colimits, the normal closure of h in X is given by the kernel of the cokernel of h. In an ideal-determined unital category, the Huq commutator is the normal closure of the Higgins commutator [Mantovani, Metere, 2010, Proposition 5.7]. We will prove this result in a normal unital category with finite colimits.

Let’s first recall the following:

2.1. Lemma. Let ℂ be a pointed category with finite colimits. For each commutative diagram as below

A ; B ; C ; X ; Y ; Z ; (1) ; \begin{matrix}k \\ g \\ f \\ s \\ r \end{matrix} ; \begin{matrix}coker(f) \\ e \end{matrix}

where e is an epimorphism, the diagram (1) is a pushout if g is the cokernel of k.

2.2. Proposition. In a normal unital category ℂ with finite colimits, the Huq commutator is the normal closure of the Higgins commutator.

Proof. Let h : H ↣ X and k : K ↣ X be a pair of subobjects of X in a normal unital category ℂ with finite colimits. Consider the commutative diagram

[H, K]_{H} X Q H ⋄ K H + K H × K (1) [h, k] φ \begin{matrix}m \\ κ_{H,K} \end{matrix} e[ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] coker(m)

in which the morphism φ is given by the universal property of cokernels. According to Lemma 2.1 the diagram (1) is a pushout. Therefore the kernel of coker(m), which, as explained above is the Huq commutator [H, K]_{Q}, is the normal closure of the Higgins commutator [H, K]_{H}.

3. Preservation of the Huq commutator by the normal closure of the regular image.

In a unital category ℂ with finite coproducts it can be seen that a pair of coterminal morphisms f : X ⟶ Z and g : Y ⟶ Z commutes if and only if there exists a (necessarily unique) morphism φ making the diagram

X × Y X + Y Z φ [f, g] [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ]

commute.

3.1. Lemma. For a pair of regular epimorphisms r : R ↠ X and s : S ↠ Y in a regular unital category ℂ with finite coproducts, the square

R × S X × Y R + S X + Y \begin{matrix}[ & 1 0 0 1 & ] \\ [ & 1 0 0 1 & ] \end{matrix} \begin{matrix}r + s \\ r × s \end{matrix}

is a pushout.

Proof. If u : R × S ⟶ T and v : X + Y ⟶ T are a pair of morphisms making the outer diagram

R × S X × Y R + S X + Y T r × s r + s [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] \begin{matrix}v = [v_{1}, v_{2}] \\ u \end{matrix} α

commute, then by precomposing with the coproduct inclusions into R + S, it can be seen that the composites v_{1} ∘ r and v_{2} ∘ s commute with cooperator u. Since r and s are regular epimorphisms and the composites v_{1} ∘ r and v_{2} ∘ s commute, it follows (see e.g [Borceux, Bourn, 2004, Proposition 1.6.4]) that the morphisms v_{1} and v_{2} also commute. Therefore, as said above, there is a unique morphism α : X × Y ⟶ T such that

α ∘ [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] = [v_{1}, v_{2}].

Moreover, since the canonical morphism from R + S to R × S is a (regular) epimorphism it can be seen that the lower triangle commutes, and clearly α is the unique morphism making the two triangles commute.

Before we state the main result of this section, let’s observe the following:

Given a morphism f : X ⟶ Y and a normal epimorphism q : X ↠ Q in a normal category ℂ with finite colimits, if t ∘ r is the regular epimorphism-monomorphism factorization of f ∘ ker(q) where ker(q) is the kernel of q, then, applying Lemma 2.1 to the commutative diagram

f(Ker(q)) ; Y ; Coker(t) ; Ker(q) ; X ; Q ; (1) ; f ; \begin{matrix}coker(t) \\ q \\ r \end{matrix} ; α ; \begin{matrix}ker(q) \\ t \end{matrix}

in which the morphism α is given by the universal property of cokernels, it follows that the diagram (1) is a pushout.

3.2. Theorem. Let h : H ↣ X and k : K ↣ X be a pair of subobjects of X and f : X ⟶ Y a morphism in a normal unital category ℂ with finite colimits. Then,

[f(H), f(K)]_{Q} = \overline{f([H, K]_{Q})}.

Proof. Let h′ ∘ u and k′ ∘ v be the regular epimorphism-monomorphism factorizations of f ∘ h and f ∘ k respectively. Consider the diagram

Q Coker(t) Y X H + K H × K f(H) + f(K) f(H) × f(K) ⌜ ⌜ φ f [h′, k′] [h, k] u × v u + v [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] qcoker(t) \begin{matrix}α \\ β \end{matrix}

in which the front face is a pushout. As explained above, the right hand face of the cube in which t is the regular image of [H, K]_{Q} ↣ X (the kernel of q) under f and α is the unique morphism given by the universal property of cokernels, is a pushout. Furthermore, since the left hand face of the cube is a pushout by Lemma 3.1, and the whole diagram commutes, there exists a unique morphism β making the back face of the cube commute. The composite of a pushout with another is a pushout, therefore the outer rectangle of the diagram

H × K f(H) × f(K) Coker(t) H + K f(H) + f(K) Y ⌜ u × v u + v [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] coker(t) β [h′, k′]

is a pushout, and hence the square

f(H) × f(K) Coker(t) f(H) + f(K) Y β [h′, k′] coker(t) [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ]

is a pushout. Since the subobject \overline{f([H, K]_{Q})} ↣ Y is the kernel of coker(t), it follows that [f(H), f(K)]_{Q} = \overline{f([H, K]_{Q})}.

Recall that each morphism f : X ⟶ Y in a regular category ℂ induces a Galois connection between the posets of subobjects Sub(X) and Sub(Y) of X and Y respectively, i.e. a pair of maps f_{∗} : Sub(X) ⟶ Sub(Y) and f^{∗} : Sub(Y) ⟶ Sub(X) defined (respectively) by taking regular images of subobjects of X under f and pulling back subobjects of Y along f, such that f_{∗}f^{∗}(L) ≤ L and H ≤ f^{∗}f_{∗}(H) for all subobjects h : H ↣ X and l : L ↣ Y of X and Y respectively.

3.3. Lemma. For a subobject h : H ↣ X of X and a morphism f : X ⟶ Y in a normal category ℂ with finite colimits, one has

\overline{f(H)} = \overline{f(\overline{H})}.

Proof. It is clear that \overline{f(H)} ≤ \overline{f(\overline{H})}. On the other hand, \overline{f(\overline{H})} ≤ \overline{f(H)} if f(\overline{H}) ≤ \overline{f(H)} or equivalently f_{∗}(\overline{H}) ≤ \overline{f_{∗}(H)}, since \overline{f(H)} is a normal subobject of Y and \overline{f(\overline{H})} is the smallest normal subobject of Y such that f(\overline{H}) ≤ \overline{f(\overline{H})}. By the Galois connection above, f_{∗}(H) ≤ \overline{f_{∗}(H)} implies H ≤ f^{∗}(\overline{f_{∗}(H)}). Hence, \overline{H} ≤ f^{∗}(\overline{f_{∗}(H)}) since f^{∗}(\overline{f_{∗}(H)}) is normal in X, and thus f_{∗}(\overline{H}) ≤ \overline{f_{∗}(H)} follows.

For a morphism f : X ⟶ Y and a pair of subobjects h : H ↣ X and k : K ↣ X of X in a normal unital category ℂ with finite colimits, since the Huq commutator is the normal closure of the Higgins commutator, by the previous lemma it immediately follows that \overline{f([H, K]_{H})} = \overline{f([H, K]_{Q})}. We further observe that

\begin{aligned}[f(H), f(K)]_{H} & ≤ \overline{[f(H), f(K)]}_{H} \\ & = [f(H), f(K)]_{Q} \\ & = \overline{f([H, K]_{Q})} \\ & = \overline{f(\overline{[H, K]}_{H})} \\ & = \overline{f([H, K]_{H})}. \end{aligned}  (1)

The following result seems to be known and can be deduced from [Mantovani, Metere, 2010, Lemma 5.11], however we insert a proof to make the paper more self-contained.

3.4. Lemma. For a pair of subobjects h : H ↣ X and k : K ↣ X of X and a morphism f : X ⟶ Y in an ideal-determined unital category ℂ, f([H, K]_{H}) = [f(H), f(K)]_{H}.

Proof. Let h′ ∘ u and k′ ∘ v be the regular epimorphism-monomorphism factorizations of f ∘ h and f ∘ k respectively. Consider the diagram

f(H) ⋄ f(K) f(H) + f(K) f(H) × f(K) H ⋄ K H + K H × K (i) (ii) eu + v u × v [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] [ \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ] \begin{matrix}m \\ κ_{H,K} \end{matrix}

where m ∘ e is the regular epimorphism-monomorphism factorization of (u + v) ∘ κ_{H,K}. Using Lemma 3.1 the square (ii) is a pushout. Therefore, since u + v is a regular epimorphism and ℂ is an ideal-determined unital category, it follows (see for instance [Mantovani, Metere, 2010, Lemma 4.1]) that the regular image m of the kernel κ_{H,K} under u + v is the kernel of the canonical morphism from f(H) + f(K) to f(H) × f(K). Hence f([H, K]_{H}) = [f(H), f(K)]_{H} follows by the uniqueness of regular images.

In the next remark we will see that Theorem 3.2 has a much simpler proof in an ideal-determined unital category.

3.5. Remark. From the previous lemma, f([H, K]_{H}) = [f(H), f(K)]_{H} for a morphism f : X ⟶ Y and a pair of subobjects h : H ↣ X and k : K ↣ X of X in an idealdetermined unital category ℂ. Now since the Huq commutator is the normal closure of the Higgins commutator and \overline{f([H, K]_{H})} = \overline{f(\overline{[H, K]}_{H})}, then

\begin{aligned}[f(H), f(K)]_{Q} & = \overline{[f(H), f(K)]}_{H} \\ & = \overline{f([H, K]_{H})} \\ & = \overline{f(\overline{[H, K]}_{H})} \\ & = \overline{f([H, K]_{Q})}. \end{aligned}  (2)

4. A characterization of the Huq commutator.

Let ℂ be a normal unital category with finite colimits. For a pair of subobjects h : H ↣ X and k : K ↣ X of X in ℂ, it is easy to check that [H, K]_{Q} ≤ \overline{H} ∧ \overline{K}. Furthermore, for a morphism f : X ⟶ Y in ℂ, \overline{f([H, K]_{Q})} = [f(H), f(K)]_{Q} (see Theorem 3.2); in particular, [f(H), f(K)]_{Q} ≤ \overline{f([H, K]_{Q})}. One can also check that the Huq commutator satisfies the monotonicity condition i.e. if H′ ≤ H and K′ ≤ K for subobjects h′ : H′ ↣ X, h : H ↣ X, k′ : K′ ↣ X and k : K ↣ X of X, then [H′, K′]_{Q} ≤ [H, K]_{Q}. In this section we will characterize the Huq commutator as the largest commutator defined for subobjects in a normal unital category ℂ with finite colimits, satisfying the monotonicity condition as well as the following two conditions:

A1 [H, K] ≤ \overline{H} ∧ \overline{K};

A2 [f(H), f(K)] ≤ \overline{f([H, K])}.

In contrast with the Huq commutator, we will provide an example of a commutator of subobjects in a normal unital category ℂ with finite colimits satisfying all the conditions above, in which the condition A2 does not necessarily imply [f(H), f(K)] = \overline{f([H, K])}.

4.1. Example. The Higgins commutator is monotone and satisfies A1. Moreover, as explained in equation (1) above, the Higgins commutator satisfies A2 in a normal unital category with finite colimits. Since in general the Higgins commutator is not always normal (see e.g [Cigoli, 2009, Example 5.3.9]), then [f(H), f(K)]_{H} ≤ \overline{f([H, K]_{H})} does not necessarily imply [f(H), f(K)]_{H} = \overline{f([H, K]_{H})}. For example if f is the inclusion of the alternating group A_{5} into S_{5} then for the two subgroups H = ⟨(1 2)(3 4)⟩ and K = ⟨(1 2)(4 5)⟩ of A_{5}, we have [f(H), f(K)]_{H} = ⟨(345)⟩ and \overline{f([H, K]_{H})} = A_{5}.

For a pair of objects H and K in a normal unital category ℂ with finite colimits, we will denote the Huq commutator of the coproduct inclusions i_{1} : H ⟶ H + K and i_{2} : K ⟶ H + K by [i_{1}(H), i_{2}(K)]_{Q}. The morphisms ker[1, 0] : H♭K ↣ H + K and ker[0, 1] : K♭H ↣ H+K will be the kernels of the induced morphisms [1, 0] : H+K ⟶ H and [0, 1] : H +K ⟶ K respectively. It is easy to check that the co-smash product H ⋄K of H and K is the meet of K♭H and H♭K (see for instance [Mantovani, Metere, 2010]). In the next lemma we will observe that the co-smash product of H and K coincides with the Huq commutator of the coproduct inclusions i_{1} : H ⟶ H +K and i_{2} : K ⟶ H +K.

4.2. Lemma. Let H and K be objects in a normal unital category ℂ with finite colimits. Then H ⋄ K = [i_{1}(H), i_{2}(K)]_{Q}.

Proof. Since the square

H × K H × K H + K H + K \begin{matrix}[ & 1 0 0 1 & ] \\ [ & 1 0 0 1 & ] \end{matrix}

is a pushout, it follows that the kernel κ_{H,K} : H ⋄ K ↣ H + K of the canonical morphism from H+K to H×K, is the Huq commutator of the coproduct inclusions i_{1} : H ⟶ H+K and i_{2} : K ⟶ H + K.

4.3. Lemma. Let h : H ↣ X and k : K ↣ X be a pair of subobjects of X in a normal unital category ℂ with finite colimits. Then

[K♭H, H♭K]_{Q} = [i_{1}(H), i_{2}(K)]_{Q}.

Proof. We see that [K♭H, H♭K]_{Q} ≤ (K♭H) ∧ (H♭K) = H ⋄ K = [i_{1}(H), i_{2}(K)]_{Q}. On the other hand, since H ≤ K♭H, K ≤ H♭K and the Huq commutator is monotone, it follows that [i_{1}(H), i_{2}(K)]_{Q} ≤ [K♭H, H♭K]_{Q}. Hence, [K♭H, H♭K]_{Q} = [i_{1}(H), i_{2}(K)]_{Q}.

4.4. Proposition. For each pair of subobjects h : H ↣ X and k : K ↣ X of X in a normal unital category ℂ with finite colimits,

[[h, k](K♭H), [h, k](H♭K)]_{Q} = [H, K]_{Q}.

Proof. Applying Theorem 3.2 and then the previous lemma, we get

\begin{aligned}[[h, k](K♭H), [h, k](H♭K)]_{Q} & = \overline{[h, k]([K♭H, H♭K]_{Q})} \\ & = \overline{[h, k]([i_{1}(H), i_{2}(K)]_{Q})} \\ & = [H, K]_{Q}. \end{aligned}  (3)

Note that the last equality holds since [h, k]([i_{1}(H), i_{2}(K)]_{Q}) is the Higgins commutator [H, K]_{H}, and in a normal unital category with finite colimits the Huq commutator is the normal closure of the Higgins commutator (see Proposition 2.2).

For a pair of subobjects h : H ↣ X and k : K ↣ X of X in a normal category ℂ with finite coproducts, since in particular H ≤ K♭H in H + K, we see that the diagram

\begin{matrix}K♭H \\ H + K \\ [h, k](K♭H) \\ H \\ X \\ [h, k] \\ i_{1} \\ h \end{matrix}

commutes and this implies that H ≤ [h, k](K♭H) in X. Similarly, K ≤ [h, k](H♭K) in X.

Now we are ready to state our main result.

4.5. Theorem. In a normal unital category ℂ with finite colimits, the Huq commutator is the largest monotone commutator defined for subobjects satisfying the conditions A1 and A2.

Proof. Suppose there is a monotone commutator [|−, −|] defined for subobjects, satisfying the conditions A1 and A2. Let h : H ↣ X and k : K ↣ X be a pair of subobjects of X. Since [|K♭H, H♭K|] ≤ K♭H ∧ H♭K and K♭H ∧ H♭K = H ⋄ K = [K♭H, H♭K]_{Q}, then \overline{[h, k]([|K♭H, H♭K|])} ≤ \overline{[h, k]([K♭H, H♭K]_{Q})}. Since H ≤ [h, k](K♭H) and K ≤ [h, k](H♭K), applying the monotonicity condition, the condition A2, Theorem 3.2 and Proposition 4.4, we get

\begin{aligned}[|H, K|] & ≤ [|[h, k](K♭H), [h, k](H♭K)|] \\ & ≤ \overline{[h, k]([|K♭H, H♭K|])} \\ & ≤ \overline{[h, k]([K♭H, H♭K]_{Q})} \\ & = [[h, k](K♭H), [h, k](H♭K)]_{Q} \\ & = [H, K]_{Q}. \end{aligned}  (4)

References

F. Borceux and D. Bourn, Mal’cev, protomodular, homological and semi-abelian categories. Math. Appl. 566, Kluwer Acad Publ., 2004.

D. Bourn, Commutator theory in regular Mal’cev categories, in Hopf Algebras and Semiabelian Categories. G. Janelidze, B. Pareigis, W. Tholen eds., the Fields Institute Commun. 43, Amer. Math. Soc., 61-75, 2004.

D. Bourn, Intrinsic centrality and associated classifying properties. Journal of Algebra 256, 126-145, 2002.

D. Bourn, Mal’cev categories and fibrations of pointed objects. Appl. Categ. Structure. 4, 307-327, 1996.

A. S. Cigoli, Centrality via internal actions and action accessibility via centralizers. Ph.D. thesis, Università degli Studi di Milano, 2009.

J. Hagemann and C. Herrmann, A concrete ideal multiplication for algebraic systems and its relation to congruence distributivity. Arch. Math. 32, 234-245, 1979.

P. J. Higgins, Groups with multiple operators. Proc. Lond. Math. Soc. 6, 366-416, 1956.

S.A Huq, Commutator, nilpotency and solvability in categories. Quart. J. Math. Oxford Ser. 19, 363-389, 1968.

G. Janelidze, L. Márki, W.Tholen and A. Ursini, Ideal-determined categories. Cah. Topol. Géom. Différ. Catég. 51, 115-125, 2010.

Z. Janelidze, The pointed subobject functor, 3 × 3 lemmas, and subtractivity of spans. Theory and Application of categories 23, 221-242, 2010.

S. Mantovani and G. Metere, Normalities and commutators. Journal of Algebra 324, 2568-2588, 2010.

M.C. Pedicchio, A categorical approach to commutator theory. Journal of Algebra 177, 647-657, 1995.

J.D.H. Smith, Mal’cev varieties. Springer L.N. in Math. 554, 1976.

Department of Mathematics, University of Namibia, 340 Mandume Ndemufayo Avenue, Pionierspark, Windhoek, Namibia

Email: vshaumbwa@unam.na
This article may be accessed at http://www.tac.mta.ca/tac/

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available from the journal’s server at http://www.tac.mta.ca/tac/. It is archived electronically and in printed paper format.

Subscription information Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors The typesetting language of the journal is TEX, and L^{A}TEX2e is required. Articles in PDF format may be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors.

Clemens Berger, Université de Nice-Sophia Antipolis: cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown(at)btinternet.com Valeria de Paiva: Nuance Communications Inc: valeria.depaiva@gmail.com Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Kathryn Hess, Ecole Polytechnique Fédérale de Lausanne: kathryn.hess@epfl.ch Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, Macquarie University: steve.lack@mq.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@buffalo.edu Tom Leinster, University of Edinburgh: Tom.Leinster@ed.ac.uk Ieke Moerdijk, Utrecht University: i.moerdijk@uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Giuseppe Rosolini, Università di Genova: rosolini@disi.unige.it Alex Simpson, University of Ljubljana: Alex.Simpson@fmf.uni-lj.si James Stasheff, University of North Carolina: jds@math.upenn.edu Ross Street, Macquarie University: ross.street@mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca